Twice the smallest of three consecutive odd integers is eleven more than the largest. Find the integers. (Enter your answers as a comma-separated list.)


Answer by CubeyThePenguin(3113)

consecutive odd integers: (x-2), x, (x+2)

2(x-2) = 11 + (x+2)
2x-4 = x+13
x = 17

integers: 15, 17, 19